how do you find the slope of the straight line that passes
through (−1, 14) and (3, 4)

Answer by stanbon(75887)

find the slope of the straight line that passes
through (−1, 14) and (3, 4)
----------------------
slope = change in y/change in x
---
slope = (14-4)/(-1-3) = 10/-4 = -5/2
========================================
cheers,
Stan H.